The table below shows hypothetical values, in billions of dollars, of different forms of money a. Use the table to calculate the M1 and M2 money supplies for each year, as well as the growth rates of the M1 and M2 money supplies from the previous year. b. Why are the growth rates of M1 and M2 so different? Explain. Currency Money market mutual fund shares Saving account deposits Money market deposit accounts Demand and checkable deposits Small denomination time deposits Travelers checks 3-month treasury bills 2015 900 680 5,500 1,214 1,000 830 4 1,986 2016 920 681 5,780 1,245 972 861 4 2,374 2017 925 679 5,968 1,274 980 1,123 2018 931 688 6,105 1,329 993 1,566 A. E. 2,436 2,502 c. Go to the St. Louis Feds data web site called FRED. Graph M1 and M2 growth on the same graph over the time period 1960 - 2017 using monthly seasonally adjusted data. Include the graph in your submission. *get help how to use the website MOV

Please do not hand write or abbreviate any of the written statements, please. It makes it very difficult to decipher. Thank you!
Answer #1

A) The M1 money supply is the sum of the rows A, E and G respectively for each year. i.e.

Total M1 money stock

2015: 900+1000+4= 1904   similarly for

2016: 920+ 972+4=1896

2017: 925+980+3= 1908

2018: 931+993+2=1926

The M2 money supply is the sum of all components i.e. from A to G for each year.

Note: 3-months treasury bills are not considered of M1 and M2 money supply.

Total M2 money stock

2015: 900+680+5500+1214+1000+830+4= 10128

2016: 920+681+5780+1245+972+861+4= 10463

2017: 925+679+5968+1274+980+1123+3= 10952

2018: 931+688+6105+1329+993+1566+2= 11614

Growth Rate= (difference between years / earlier years) * 100

M1 growth rate 2015 to 2016:

(1896-1904) / 1904*100= (-8/1904)*100 = -0.42%

M1 growth rate 2016 to 2017:

(1908-1896)/1896*100 = 12/1896*100= 0.63%

M1 growth rate 2017 to 2018:

(1926-1908)/1908*100 = 18/1908*100 = 0.94%

M2 growth rate 2015 to 2016:

(10463-10128)/10128*100 = 335/10128*100 = 3.3%

M2 growth rate 2016 to 2017:

(10952-10463)/10463*100 = 489/10463*100 = 4.6%

M2 growth rate 2017 to 2018:

(11614-10952)/10952*100 = 662/10952*100 = 6.04%

B) The growth rate of M1 and M2 money supply are so different because M1 is the narrow measure of money and includes less components into it where as M2 is a broad measures of money supply and it includes M1 components of money supply as well as other components too.